A particle moves along straight line works at a time t a distance x of particle from is given by x is equal to 40 + 12 - 3 Cube how long would the particle travel before coming to rest

Grade:11

1 Answers

Arun
5 years ago
 
We have equation x= 40+12t-t^3. So initially take t=0  => x=40m.
Now differtiate the equation with respect to time(t) which is equal to velocity (v) {v=dx/dt} now put v=0 as the particle is stopping you will get value of t. Put this value in "40+12t-t^3" and you will get final distance.on Subtracting the initial distance from this final distance we'll get x=16m.
 
Regards
Arun (askIITians forum expert)